OpenStudy (anonymous):

let V be the space spanned by v1=cos^2x ,v2=sin^2x ,v3=cos2x a)show that S={ v1,v2,v3} is not a basis for V b) find a basis for V

OpenStudy (anonymous):

What's a basis?

OpenStudy (abb0t):

\(cos(2x) = cos^2x - sin^2x\)

OpenStudy (anonymous):

:----- A basis of a vector space V is a linearly independent spanning set. Since it's already been given that {v1,v2,v3} spans V You want to show that v1,v2, and v3 are linearly dependent.

OpenStudy (abb0t):

Thus the three functions are not linearly independent and do not form a minimal spanning set (basis)

OpenStudy (abb0t):

A basis is cos 2x because \(cos (2x) = 1-2sin^2(x)\) and \(cos(2x) = 2cos^2(x) -1\) i.e. \(sin^2(x) \) and \(cos^2(x) \)can be written in terms of cos(2x) Similarly you can use \(sin^2(x)\) or \(cos^2(x)\) as a basis. Any ONE of these can be used as a basis.

OpenStudy (anonymous):

Wrong, @abb0t It's not sufficient that you are able to express \(\cos^2 x\) in terms of \(\cos 2x\) it must be a LINEAR combination, which it is not. @yantinearsita yes, you may use any TWO of these as a basis. I'd stick to \(\cos^2 x \) and \(\sin^2 x\) though.

OpenStudy (anonymous):

oke, thank you so much @PeterPan and @abb0t for helping me to solve my homework

OpenStudy (anonymous):

Remember, though, one of them is not enough for a basis...

OpenStudy (anonymous):

ok, i get it